Question 16(1 point) height in cm weight in k 1.96r
Answer #1

Question 16:
d, High emetic risk
This medicine combination for 90% high frequency of emesis. this medication 5HT3, NK1 and corticosteroids appear to be interchangeable.
Question 17:
c, 1.96m2
Height in cm = 165cm
weight in kg = 84 kg
BSA = Height in cm X weight in kg / 3600
165 X 84/ 3600 = 13860 / 3600 = 3.85
to change in meter =3.85 divided in to 2 = 1.96m2